Operations Report – January 6th

SOL: 8

Name of person filing report: Kasey Hilton

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Generator (hours run): 17hr 58min; Turned on last night (05Jan2019) at 15:05; Turned off this morning (06Jan2019) at 9:03; Turned on tonight (06Jan2019) at 13:24

Solar SOC – Turned on (05Jan2019) 80%; Turned off (06Jan2019) 80%; Turned on (06Jan2019) 72%

Diesel Reading – 73%

Propane Reading – 43%

Ethanol Free Gasoline – Not in use

Water (auxiliary tank) – Not in use

Water (static tank) – About 70%; 345 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: Not applicable

Water in GreenHab – About 70%; 215 gallons (assuming that the tank is 300 gallons)

Water (loft) – At level marker 12

Static to Loft Pump used – Yes; At 19:33 to refill tank

Water Meter: 01398548

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Yes

Hours: 65.5

Beginning charge: 100%

Ending charge: 55%

Currently charging: Yes

Opportunity rover used: No; Still out of brake fluid

Hours: 45.4

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 63.7

Beginning charge: 100%

Ending charge: 71%

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: None (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: Not applicable

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: None

Notes on ATVs: All ATVs were started and run to prevent the batteries from freezing

HabCar used and why, where? Not used

CrewCar used and why, where? Off site

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Batteries in suits #7 and #8 were replaced since each of their batteries had a voltage less than 12V

Summary of Hab operations: EVA airlock light fixed

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report